# System Password Security: Attack and Defense Mechanisms

### v1

@@ 1. Introduction
In the era of the digital economy, information systems are deeply integrated into social production and daily life. As the first line of defence in access control, password authentication remains the most common authentication method due to its simple implementation and low cost. However, with the continuous evolution of attack techniques, password security faces unprecedented challenges. The computational power of high-performance brute-force cracking tools increases rapidly, large-scale dictionary libraries expand constantly, and the success rate of social-engineering attacks combined with machine learning rises significantly, so the traditional defence system undergoes a severe test (sen2012applied, ). Therefore, the study of system password cracking attacks and defence techniques is of great theoretical and practical significance. This research not only enriches the network security system at the academic level, but also provides direct guidance for practice. The systematic exploration of password attack mechanisms and defence mechanisms helps design a more secure password strategy and storage scheme, and enhances the public’s security awareness and risk perception, thus promoting the construction of an overall network security culture. At the same time, effective password defence technology applies to all kinds of information systems to reduce the risk of data leakage and business interruption, protect personal privacy and corporate assets, and strengthen the network security protection capability at the national level (bybee1986science, ). Based on the above background, this paper focuses on the typical attack methods of system password cracking and their defense strategies. Firstly, we review the research progress in the field of password security at home and abroad, and analyze the common methods and their technical characteristics, such as brute-force cracking, dictionary attacks, and rainbow table attacks, etc. (gong2025information, ). Then, we design experiments, use John the Ripper and Hashcat tools to test the hash value of passwords generated by different hash algorithms (MD5, SHA-256, bcrypt) with brute-force and dictionary cracking, and evaluate the effectiveness of password complexity policy, salt, and slow hash algorithms. Finally, we discuss new defense schemes, such as multi-factor authentication, risk-adaptive authentication, and honeywords, along with their optimization ideas in light of the latest research results (dell2010password, ).
@@ 2.1. MD5
The overall process of the MD5 algorithm is illustrated in Fig. 1. MD5 (Message-Digest Algorithm 5) is a widely used hash function that produces a 128-bit output.
@@
The basic procedure is as follows. The input message is first padded by appending a single ”1” bit and enough ”0” bits so that the total length is congruent to 448 modulo 512, and then a 64-bit little-endian representation of the original message length is appended. After padding, the message is divided into multiple 512-bit blocks, each treated as 16 little-endian 32-bit words (bonneau2010password, ). The algorithm initializes four 32-bit registers $A$, $B$, $C$, and $D$ with the values
@@
For each data block, MD5 performs four rounds of nonlinear iterative operations with a total of 64 steps. Each round uses a distinct Boolean function $F$, $G$, $H$, or $I$, and a constant $T[i]$ derived from the sine function. Register updates follow the predefined order and left-rotation parameter $s$ (lu2025movescanner, ).
@@
where $(A,B,C,D)\rightarrow(D,A,B,C)$ after each step. After all data blocks are processed, the final values of $A$, $B$, $C$, and $D$ are concatenated in little-endian order to produce the 128-bit hash value. Although MD5 is computationally efficient, it is proven vulnerable to collision attacks and is no longer recommended for password storage or other security-critical applications. More secure hash algorithms, such as SHA-256 and bcrypt, are now preferred (bowen2009baiting, ).
@@ 2.2. SHA-256
SHA-256, as a cryptographic secure hash algorithm, belongs to the SHA-2 family. During the hashing process, the input data is first padded, with the original message length appended in big-endian order. The padded data is divided into 512-bit blocks, and each block is processed through 64 rounds of a compression function. In each round, different logical functions such as bitwise operations and modular addition, as well as predefined constants, are used to iteratively update eight 32-bit intermediate hash values (kong2025uechecker, ). Finally, these eight intermediate values are concatenated to form a 256-bit hash. The core security of SHA-256 relies on its complex nonlinear structure and collision resistance. Although no effective practical collision attack has yet been found, potential threats from quantum computing must still be considered (cohen2006use, ). The SHA-256 algorithm process is shown in Fig. 2.
@@
(1) Message Preprocessing: An input message $M$ of arbitrary length is padded and divided into $N$ blocks of 512 bits ($M_{0}$ to $M_{N-1}$). Padding requires the message length modulo $512$ to equal $448$, so a single bit ”1” is appended, followed by enough ”0” bits until the length reaches $448\pmod{512}$. Finally, an $8$-byte big-endian representation of the original message length is appended.
(2) Hash Initialization: Eight 32-bit registers AAHH are initialized as $A=0\text{x}6a09e667,\quad B=0\text{x}bb67ae85,\quad C=0\text{x}3c6ef372,\quad D=0\text{x}a54ff53a,E=0\text{x}510e527f,\quad F=0\text{x}9b05688c,\quad G=0\text{x}1f83d9ab,\quad H=0\text{x}5be0cd19.$ These constants are derived from the fractional parts of the square roots of the first eight prime numbers (wang2019birthday, ).
@@
where $\mathbin{\text{ROR}}$ denotes cyclic right rotation and $\mathbin{\text{SHR}}$ denotes logical right shift.
(4) Compression Function: For each of the 64 rounds, the following Boolean functions are applied
@@
with $K_{t}$ denoting the $t$-th constant derived from the cube roots of the first 64 prime numbers. The registers are updated according to $H_{t+1}=G_{t},G_{t+1}=F_{t},F_{t+1}=E_{t},E_{t+1}=D_{t}+T1_{t},D_{t+1}=C_{t},C_{t+1}=B_{t},B_{t+1}=A_{t},A_{t+1}=T1_{t}+T2_{t}.$
@@ 2.3. Bcrypt algorithm
Bcrypt is an adaptive hash algorithm specifically designed for password storage. Its core idea is to enhance resistance to brute-force and rainbow table attacks by introducing random salts, configurable computational cost, and memory-intensive operations. The bcrypt algorithm process is illustrated in the figure. The final output is a 60-character string that includes the algorithm identifier, the cost value, a 22-character salt, and a 31-character hash. The security advantages of bcrypt can be summarized in three aspects (yegireddi2016survey, ). First, the computational cost is adjustable because the cost parameter determines the number of iterations $(2^{\text{cost}})$, which enables the computation to scale with hardware performance. Second, a random salt is mandatory since a 16-byte (128-bit) cryptographically secure random salt is automatically generated for each encryption, which prevents rainbow table attacks. Third, the computation is memory-intensive because the key expansion process requires substantial memory access, which helps resist GPU or ASIC acceleration (zhang2025penetration, ). The bcrypt algorithm process is shown in Fig. 3.
@@
(1) Input and Initialization. The user password is first encoded into a UTF-8 byte sequence $M$ and truncated to 72 bytes. The iteration parameter cost is set with a default value of 10, corresponding to $2^{10}$ iterations, and a 16-byte random salt is generated.
(2) Key Expansion. The P-array contains 18 subkeys of 32 bits each, initialized from the hexadecimal expansion of $\pi$. The S-box consists of four lookup tables of size $256\times 32$ bits, also initialized from $\pi$. The password sequence $M$ is divided into 32-bit blocks $W_{0}\ldots W_{m-1}$ in big-endian order, and cyclic padding is applied if necessary (peng2025multicfv, ). The password and salt are alternately used to update the P-array and S-box through the EksBlowfishSetup algorithm
@@
Afterward, the Blowfish encryption function is repeatedly applied to mix in the salt $S$, and this process is repeated $2^{\text{cost}}$ times according to the cost parameter (jo2015new, ).
(3) Data Encryption. The fixed plaintext string ”OrpheanBeholderScryDoubt” (24 bytes, three 64-bit blocks) is encrypted in CTR mode using 64 rounds of Blowfish. Each round consists of three steps. First, the P-array permutation is performed by XOR with the current subkey. Second, S-box substitution is carried out through non-linear table lookup and modular addition. Third, the Feistel structure is updated according to
@@
(4) Output Format. The three encrypted blocks are concatenated to produce a 24-byte ciphertext (zou2025malicious, ). The final output string is represented as
@@ 2.4. Password Cracking Attack
In the field of information security, password cracking attacks evolve into multiple mature technical lines, each with distinct principles and application scenarios. With the advancement of computational power and the evolution of attack methods, modern password cracking exhibits characteristics of intelligence, automation, and scalability. Attackers typically select the most effective single method according to the protection level of the target system or combine multiple methods to increase success probability (zhang2025risk, ). Given the widespread deployment of online protections such as account lockout and human verification, attackers more commonly perform offline comparisons of stored password hashes to recover plaintext passwords. The following highlights several typical hash-targeted attack techniques (karuna2018generating, ).
(1) Brute-force attack is the most primitive yet deterministic approach. Its core idea is to enumerate all possible character combinations to find a plaintext that matches the target hash. Although it is extremely time-consuming for high-complexity passwords, the method is universal and applicable to any password type. Modern brute-force tools optimize search strategies through layered search that prioritizes high-frequency character combinations. They also leverage GPU parallelism or distributed computing to substantially increase search speed (shen2025blockchain, ).
(2) Dictionary attack exploits statistical regularities and user habits. Attackers construct dictionaries containing common passwords, usernames, phrases, and their variants, such as case substitutions or numeric and symbol suffixes. These entries are compared against the target hash. The effectiveness of the attack relies heavily on the quality of the dictionary and the rule engine. Advanced tools, such as those based on Markov chains or probabilistic models, infer user preferences from leaked data. This enables the cracking of many common passwords at a fraction of the computational cost of brute force in practice (temoshok2022digital, ).
(3) Rainbow table attack applies a time–memory trade-off by precomputing and storing large mappings between hashes and candidate plaintexts using chain structures. Reduction functions map hash values back to candidate plaintexts, allowing attackers to trace a matching hash along a chain to potential original passwords. Compared to pure enumeration, rainbow tables reduce lookup time at the expense of storage overhead. Their effectiveness is significantly weakened when salts are incorporated into hash storage (wang2025ai, ).
@@ 2.5. Multi-Factor Authentication and Account Lock Mechanism
Multi-factor authentication (MFA) requires users to provide two or more different types of verification factors when logging in. These factors are generally classified into three categories, including knowledge factors such as passwords, possession factors such as mobile devices and smart cards, and biometric factors such as fingerprints and facial recognition. By combining independent elements, MFA adds a strong layer of defense so that an attacker who compromises one factor still cannot gain account access. Modern MFA implementations include time-based one-time passwords (TOTP), push-notification approvals, and biometric verification. Emerging passwordless technologies, such as the FIDO2 standard, integrate public-key cryptography with biometrics, improving both security and user experience. From a system-architecture perspective, advanced MFA solutions often adopt risk-based adaptive authentication, dynamically adjusting requirements according to contextual signals such as device fingerprints, geographic location, and behavioral patterns (omolara2018novel, ).
The account lockout mechanism is an essential security measure designed to prevent unauthorized access and malicious attacks and represents one of the most effective countermeasures against brute-force attempts. Its principle is straightforward, as the system detects multiple consecutive incorrect password entries, abnormal login attempts, or other security risks and automatically locks the account, thereby blocking further login attempts. This mechanism makes it nearly impossible for attackers to compromise an account through brute-force or dictionary attacks, effectively safeguarding the confidentiality and integrity of user data and assets (liu2025empirical, ). Account lockout is triggered automatically by the system or manually by an administrator. Once locked, a user must either wait for an automatic unlock after a specified period or perform additional identity verification to regain access. The process can also incorporate challenge-response mechanisms such as CAPTCHA to distinguish legitimate users from automated attack tools. Although highly effective, improper configuration of the lockout mechanism may be exploited by attackers to launch denial-of-service (DoS) attacks. To address this risk, modern account-protection systems detect distributed brute-force attacks and accurately identify and block malicious activity even when attackers use varying IP addresses and user agents (rao2006data, ).
@@ 2.6. Salting Treatment and Risk-Adaptive Authentication
Salted password hashing is a critical security technique for protecting stored credentials. Its core principle is to introduce a globally unique random salt for each user password, ensuring that even if different accounts share the same password, the resulting hash outputs are entirely distinct. Consequently, an attacker attempting a rainbow-table attack must construct a separate table for each user, fundamentally undermining the feasibility of such attacks. In practical deployment, combining salting with key derivation functions that resist brute force, such as PBKDF2, bcrypt, or Argon2, forces attackers to expend substantial time and computational resources even when equipped with high-performance hardware (schechter2009s, ). The security of salted password hashing depends on three key requirements. First, the salt must be unpredictable and unique for each user. Second, algorithm cost parameters, including iteration count and memory requirement, must be chosen to increase the computational cost of attacks. Third, protective measures must be comprehensively integrated across storage, transmission, and verification. These measures jointly build a multi-layered defense against credential compromise. Numerous international standards bodies, including NIST, OWASP, and ISO/IEC, provide explicit guidance on salting, highlighting its critical role in password security (ross2005stronger, ).
Risk-adaptive authentication is an intelligent mechanism that dynamically adjusts authentication requirements based on real-time risk assessment. The system evaluates multiple contextual signals, including login time, geographic location, device fingerprint, and network environment, to generate a risk score for each login attempt. Low-risk requests require only password verification. Medium-risk requests require additional MFA. High-risk requests result in the login being blocked. During implementation, excessive blocking of legitimate users must be avoided. Providing an appeal or manual review channel is recommended to prevent unintended account lockouts (peng2025mining, ). The overall operating principle of the risk-adaptive authentication system is illustrated in Fig. 4.

### v2

@@ 1. Introduction
In the era of the digital economy, information systems are deeply integrated into social production and daily life. As the first line of defence in access control, password authentication remains the most common authentication method due to its simple implementation and low cost(Byun et al., 2024). However, with the continuous evolution of attack techniques, password security faces unprecedented challenges. The computational power of high-performance brute-force cracking tools increases rapidly, large-scale dictionary libraries expand constantly, and the success rate of social-engineering attacks combined with machine learning rises significantly, so the traditional defence system undergoes a severe test (Sen, 2012). Therefore, the study of system password cracking attacks and defence techniques is of great theoretical and practical significance. This research not only enriches the network security system at the academic level, but also provides direct guidance for practice(Touil et al., 2024). The systematic exploration of password attack mechanisms and defence mechanisms helps design a more secure password strategy and storage scheme, and enhances the public’s security awareness and risk perception, thus promoting the construction of an overall network security culture(He et al., 2024). At the same time, effective password defence technology applies to all kinds of information systems to reduce the risk of data leakage and business interruption, protect personal privacy and corporate assets, and strengthen the network security protection capability at the national level (Bybee and Mau, 1986). Based on the above background, this paper focuses on the typical attack methods of system password cracking and their defense strategies(Saadi et al., 2024). Firstly, we review the research progress in the field of password security at home and abroad, and analyze the common methods and their technical characteristics, such as brute-force cracking, dictionary attacks, and rainbow table attacks, etc. (Gong et al., 2025). Then, we design experiments, use John the Ripper and Hashcat tools to test the hash value of passwords generated by different hash algorithms (MD5, SHA-256, bcrypt) with brute-force and dictionary cracking, and evaluate the effectiveness of password complexity policy, salt, and slow hash algorithms(He et al., 2024). Finally, we discuss new defense schemes, such as multi-factor authentication, risk-adaptive authentication, and honeywords, along with their optimization ideas in light of the latest research results (Dell’Amico et al., 2010).
@@ 2.1. MD5
The overall process of the MD5 algorithm is illustrated in Fig. 1. MD5 (Message-Digest Algorithm 5) is a widely used hash function that produces a 128-bit output(Alzahrani, 2024).
@@
The basic procedure is as follows. The input message is first padded by appending a single ”1” bit and enough ”0” bits so that the total length is congruent to 448 modulo 512, and then a 64-bit little-endian representation of the original message length is appended(Almasoud et al., 2024). After padding, the message is divided into multiple 512-bit blocks, each treated as 16 little-endian 32-bit words (Bonneau and Preibusch, 2010). The algorithm initializes four 32-bit registers $A$, $B$, $C$, and $D$ with the values
@@
For each data block, MD5 performs four rounds of nonlinear iterative operations with a total of 64 steps. Each round uses a distinct Boolean function $F$, $G$, $H$, or $I$, and a constant $T[i]$ derived from the sine function. Register updates follow the predefined order and left-rotation parameter $s$ (Luo et al., 2025).
@@
where $(A,B,C,D)\rightarrow(D,A,B,C)$ after each step. After all data blocks are processed, the final values of $A$, $B$, $C$, and $D$ are concatenated in little-endian order to produce the 128-bit hash value(Zaikin, 2024). Although MD5 is computationally efficient, it is proven vulnerable to collision attacks and is no longer recommended for password storage or other security-critical applications. More secure hash algorithms, such as SHA-256 and bcrypt, are now preferred (Bowen et al., 2009).
@@ 2.2. SHA-256
SHA-256, as a cryptographic secure hash algorithm, belongs to the SHA-2 family. During the hashing process, the input data is first padded, with the original message length appended in big-endian order(Alimoglu et al., 2025). The padded data is divided into 512-bit blocks, and each block is processed through 64 rounds of a compression function. In each round, different logical functions such as bitwise operations and modular addition, as well as predefined constants, are used to iteratively update eight 32-bit intermediate hash values (Kong et al., 2025). Finally, these eight intermediate values are concatenated to form a 256-bit hash. The core security of SHA-256 relies on its complex nonlinear structure and collision resistance(Li et al., 2024). Although no effective practical collision attack has yet been found, potential threats from quantum computing must still be considered (Cohen, 2006). The SHA-256 algorithm process is shown in Fig. 2.
@@
(1) Message Preprocessing: An input message $M$ of arbitrary length is padded and divided into $N$ blocks of 512 bits ($M_{0}$ to $M_{N-1}$). Padding requires the message length modulo $512$ to equal $448$, so a single bit ”1” is appended, followed by enough ”0” bits until the length reaches $448\pmod{512}$. Finally, an $8$-byte big-endian representation of the original message length is appended(Wang and Li, 2025).
(2) Hash Initialization: Eight 32-bit registers AAHH are initialized as $A=0\text{x}6a09e667,\quad B=0\text{x}bb67ae85,\quad C=0\text{x}3c6ef372,\quad D=0\text{x}a54ff53a,E=0\text{x}510e527f,\quad F=0\text{x}9b05688c,\quad G=0\text{x}1f83d9ab,\quad H=0\text{x}5be0cd19.$ These constants are derived from the fractional parts of the square roots of the first eight prime numbers (Wang et al., 2019).
@@
where $\mathbin{\text{ROR}}$ denotes cyclic right rotation and $\mathbin{\text{SHR}}$ denotes logical right shift(Salih and Kashmar, 2024).
(4) Compression Function: For each of the 64 rounds, the following Boolean functions are applied(Aydin et al., 2025)
@@
with $K_{t}$ denoting the $t$-th constant derived from the cube roots of the first 64 prime numbers(Zhou et al., 2025a). The registers are updated according to $H_{t+1}=G_{t},G_{t+1}=F_{t},F_{t+1}=E_{t},E_{t+1}=D_{t}+T1_{t},D_{t+1}=C_{t},C_{t+1}=B_{t},B_{t+1}=A_{t},A_{t+1}=T1_{t}+T2_{t}.$
@@ 2.3. Bcrypt algorithm
Bcrypt is an adaptive hash algorithm specifically designed for password storage. Its core idea is to enhance resistance to brute-force and rainbow table attacks by introducing random salts, configurable computational cost, and memory-intensive operations(Al Azhar and Widiati, 2025). The bcrypt algorithm process is illustrated in the figure. The final output is a 60-character string that includes the algorithm identifier, the cost value, a 22-character salt, and a 31-character hash. The security advantages of bcrypt can be summarized in three aspects (Yegireddi and Kumar, 2016). First, the computational cost is adjustable because the cost parameter determines the number of iterations $(2^{\text{cost}})$, which enables the computation to scale with hardware performance. Second, a random salt is mandatory since a 16-byte (128-bit) cryptographically secure random salt is automatically generated for each encryption, which prevents rainbow table attacks(Guo et al., 2024). Third, the computation is memory-intensive because the key expansion process requires substantial memory access, which helps resist GPU or ASIC acceleration (Zhang et al., 2025a). The bcrypt algorithm process is shown in Fig. 3.
@@
(1) Input and Initialization. The user password is first encoded into a UTF-8 byte sequence $M$ and truncated to 72 bytes. The iteration parameter cost is set with a default value of 10, corresponding to $2^{10}$ iterations, and a 16-byte random salt is generated(Shi et al., 2025).
(2) Key Expansion. The P-array contains 18 subkeys of 32 bits each, initialized from the hexadecimal expansion of $\pi$. The S-box consists of four lookup tables of size $256\times 32$ bits, also initialized from $\pi$. The password sequence $M$ is divided into 32-bit blocks $W_{0}\ldots W_{m-1}$ in big-endian order, and cyclic padding is applied if necessary (Peng et al., 2025b). The password and salt are alternately used to update the P-array and S-box through the EksBlowfishSetup algorithm
@@
Afterward, the Blowfish encryption function is repeatedly applied to mix in the salt $S$, and this process is repeated $2^{\text{cost}}$ times according to the cost parameter (Jo and Yoon, 2015).
(3) Data Encryption. The fixed plaintext string ”OrpheanBeholderScryDoubt” (24 bytes, three 64-bit blocks) is encrypted in CTR mode using 64 rounds of Blowfish(Castelo et al., 2024). Each round consists of three steps. First, the P-array permutation is performed by XOR with the current subkey. Second, S-box substitution is carried out through non-linear table lookup and modular addition(Haryoputranto et al., 2025). Third, the Feistel structure is updated according to
@@
(4) Output Format. The three encrypted blocks are concatenated to produce a 24-byte ciphertext (Zou et al., 2025). The final output string is represented as
@@ 2.4. Password Cracking Attack
In the field of information security, password cracking attacks evolve into multiple mature technical lines, each with distinct principles and application scenarios. With the advancement of computational power and the evolution of attack methods, modern password cracking exhibits characteristics of intelligence, automation, and scalability(Saputra et al., 2025). Attackers typically select the most effective single method according to the protection level of the target system or combine multiple methods to increase success probability (Zhang et al., 2025b). Given the widespread deployment of online protections such as account lockout and human verification, attackers more commonly perform offline comparisons of stored password hashes to recover plaintext passwords. The following highlights several typical hash-targeted attack techniques (Karuna et al., 2018).
(1) Brute-force attack is the most primitive yet deterministic approach. Its core idea is to enumerate all possible character combinations to find a plaintext that matches the target hash. Although it is extremely time-consuming for high-complexity passwords, the method is universal and applicable to any password type(Reddy et al., ). Modern brute-force tools optimize search strategies through layered search that prioritizes high-frequency character combinations. They also leverage GPU parallelism or distributed computing to substantially increase search speed (Shen et al., 2025).
(2) Dictionary attack exploits statistical regularities and user habits. Attackers construct dictionaries containing common passwords, usernames, phrases, and their variants, such as case substitutions or numeric and symbol suffixes. These entries are compared against the target hash(He et al., 2024). The effectiveness of the attack relies heavily on the quality of the dictionary and the rule engine. Advanced tools, such as those based on Markov chains or probabilistic models, infer user preferences from leaked data. This enables the cracking of many common passwords at a fraction of the computational cost of brute force in practice (Temoshok et al., 2022).
(3) Rainbow table attack applies a time–memory trade-off by precomputing and storing large mappings between hashes and candidate plaintexts using chain structures(Xu et al., 2025). Reduction functions map hash values back to candidate plaintexts, allowing attackers to trace a matching hash along a chain to potential original passwords. Compared to pure enumeration, rainbow tables reduce lookup time at the expense of storage overhead. Their effectiveness is significantly weakened when salts are incorporated into hash storage (Wang and Li, 2025).
@@ 2.5. Multi-Factor Authentication and Account Lock Mechanism
Multi-factor authentication (MFA) requires users to provide two or more different types of verification factors when logging in. These factors are generally classified into three categories, including knowledge factors such as passwords, possession factors such as mobile devices and smart cards, and biometric factors such as fingerprints and facial recognition(Tihanyi et al., 2024). By combining independent elements, MFA adds a strong layer of defense so that an attacker who compromises one factor still cannot gain account access. Modern MFA implementations include time-based one-time passwords (TOTP), push-notification approvals, and biometric verification. Emerging passwordless technologies, such as the FIDO2 standard, integrate public-key cryptography with biometrics, improving both security and user experience. From a system-architecture perspective, advanced MFA solutions often adopt risk-based adaptive authentication, dynamically adjusting requirements according to contextual signals such as device fingerprints, geographic location, and behavioral patterns (Omolara et al., 2018).
The account lockout mechanism is an essential security measure designed to prevent unauthorized access and malicious attacks and represents one of the most effective countermeasures against brute-force attempts. Its principle is straightforward, as the system detects multiple consecutive incorrect password entries, abnormal login attempts, or other security risks and automatically locks the account, thereby blocking further login attempts. This mechanism makes it nearly impossible for attackers to compromise an account through brute-force or dictionary attacks, effectively safeguarding the confidentiality and integrity of user data and assets (Liu et al., 2025). Account lockout is triggered automatically by the system or manually by an administrator. Once locked, a user must either wait for an automatic unlock after a specified period or perform additional identity verification to regain access. The process can also incorporate challenge-response mechanisms such as CAPTCHA to distinguish legitimate users from automated attack tools. Although highly effective, improper configuration of the lockout mechanism may be exploited by attackers to launch denial-of-service (DoS) attacks. To address this risk, modern account-protection systems detect distributed brute-force attacks and accurately identify and block malicious activity even when attackers use varying IP addresses and user agents (Rao, 2006).
@@ 2.6. Salting Treatment and Risk-Adaptive Authentication
Salted password hashing is a critical security technique for protecting stored credentials. Its core principle is to introduce a globally unique random salt for each user password, ensuring that even if different accounts share the same password, the resulting hash outputs are entirely distinct. Consequently, an attacker attempting a rainbow-table attack must construct a separate table for each user, fundamentally undermining the feasibility of such attacks. In practical deployment, combining salting with key derivation functions that resist brute force, such as PBKDF2, bcrypt, or Argon2, forces attackers to expend substantial time and computational resources even when equipped with high-performance hardware (Schechter et al., 2009). The security of salted password hashing depends on three key requirements. First, the salt must be unpredictable and unique for each user. Second, algorithm cost parameters, including iteration count and memory requirement, must be chosen to increase the computational cost of attacks. Third, protective measures must be comprehensively integrated across storage, transmission, and verification. These measures jointly build a multi-layered defense against credential compromise. Numerous international standards bodies, including NIST, OWASP, and ISO/IEC, provide explicit guidance on salting, highlighting its critical role in password security (Ross et al., 2005).
Risk-adaptive authentication is an intelligent mechanism that dynamically adjusts authentication requirements based on real-time risk assessment. The system evaluates multiple contextual signals, including login time, geographic location, device fingerprint, and network environment, to generate a risk score for each login attempt. Low-risk requests require only password verification. Medium-risk requests require additional MFA. High-risk requests result in the login being blocked. During implementation, excessive blocking of legitimate users must be avoided. Providing an appeal or manual review channel is recommended to prevent unintended account lockouts (Peng et al., 2025a). The overall operating principle of the risk-adaptive authentication system is illustrated in Fig. 4.